A 10,000-kg rocket blasts off from earth with a uniform upward acceleration of 2.00 m/s2 and feels no air resistance. What is the upward thrust force its engines must provide during this acceleration?

Answer by ikleyn(52787)

.

The force must be   F = m*(a+g) = 10000 kg * (2.00 + 10) m/s^2 = 120000 newtons

                (Newton' second law).


Here  g = 10 m/s^2  is the  Earth gravity acceleration,  which also should be overtaken.

Solved,  answered and explained.